OpenStudy (anonymous):

evaluate (5 • 8 – 12)(36 ÷ 9 • 4)

OpenStudy (anonymous):

det cat us cute

OpenStudy (anonymous):

Do you at least have a answer you think is right?

OpenStudy (anonymous):

thank you I just adopted her

OpenStudy (anonymous):

no need to kno cuz idont know how to evaluate I am horrible in math

OpenStudy (swolesammy):

Start with P E MD AS

OpenStudy (anonymous):

that's what I did but its still confusing

OpenStudy (anonymous):

@TorilovesVampires

OpenStudy (anonymous):

@jjamz87

OpenStudy (swolesammy):

I got 448

OpenStudy (kva1992):

(5 • 8 – 12)(36 ÷ 9 • 4) (40-12)(4*4) (28)(16) 448

OpenStudy (jjamz87):

(40-12)(16) (28)(16) 448

OpenStudy (kva1992):

yeah its 448

OpenStudy (anonymous):

got that too

OpenStudy (texaschic101):

(5 * 8 - 12) (36/9 * 4) = (40 - 12) (4 * 4) = (28)(16) = 448

OpenStudy (anonymous):

thank you all

OpenStudy (anonymous):

\[_{\sqrt[ \sqrt{\sqrt{\sqrt{ \sqrt[\sqrt[\sqrt[ \prime \prime \prime _{_{ _{ _{^{^{ \sqrt{ \sqrt{\sqrt[\prime _{^{?}}]{?}}}}}}}}^{?}}^{?}]{?}]{?}]{?}}}}]{?}}\]

OpenStudy (anonymous):

uggh sorry that was my lil bro